OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (3, (3/2) pi). (0, 3) (0, -3) (-3, 0) (3, 0)

OpenStudy (chaise):

Draw a triangle with length R=3 and angle theta=3/2*pi. Remember: x=rcos(theta) y=rsin(theta)

OpenStudy (chaise):

Make sure your calculator is in radians.

OpenStudy (anonymous):

i dont have a calculator :x

OpenStudy (anonymous):

@shifuyanli @CREEPYMLPMASTER

OpenStudy (anonymous):

sorry not good with math :/

OpenStudy (jdoe0001):

\(\large { \begin{array}{cccllll} ( 3&,&\frac{3\pi }{2})\\ r&&\theta \end{array}\qquad \begin{cases} {\color{brown}{ x}}=rcos(\theta)\\ {\color{blue}{ y}}=rsin(\theta) \end{cases}\qquad ({\color{brown}{ \square ?}},{\color{blue}{ \square ? }}) }\)

OpenStudy (anonymous):

x=3 y=idk ??

OpenStudy (jdoe0001):

well. why not use the identities should be pretty straighforward enough as chaise said make sure your calculator is in Radians mode

OpenStudy (jdoe0001):

notice what "r" is notice what the angle is notice what "x" equals and notice what "y" equals

OpenStudy (anonymous):

im sorry.. i really have no clue

OpenStudy (jdoe0001):

well have you covered polar coordinates yet?

OpenStudy (anonymous):

ehh somewhat..

OpenStudy (jdoe0001):

well... you may want to cover that material first then or brush up a little more

OpenStudy (anonymous):

okay

OpenStudy (chaise):

The conversion from polar to rectangular is: \[(r, \theta)=(rcos(\theta),rsin(\theta))\]